# A Distributional Treatment of Real2Sim2Real for Vision-Driven Deformable Linear Object Manipulation

### v1

@@ IV-B Simulation & Hardware setup
We setup our simulation in IsaacGym [31]. Our simulated environment contains a Franka Emika Panda 7-DoF robot arm equipped with a parallel gripper, which is on top of a tabletop workspace. On top of this workspace, a blue-coloured DLO exists, which is simulated using the corotational finite element method of the FleX physics engine. A small green spherical object is the target of our visuomotor control task. Across all our parameter inference and policy learning experiments, the simulated DLO is parameterised in the $[1e3,5e4]$ (Pa) range for Young’s modulus and the $[195,305]$ (mm) range for length. Thus, median of our simulated parameter space is $(2.55e4\;\text{Pa},\;250\text{mm})$.
Our real-world setup replicates the simulation setup as closely as possible. We extend an open-source sample-efficient robotic RL infrastructure [30] for the purpose of evaluating PPO policies [32]. For visual observations, we use a RealSense D435i camera capturing $60$ fps, mounted to view our workspace from the right side. To avoid the overhead of a more elaborate sim and real camera pose calibration, we select our real-world camera placement such that the captured images qualitatively approximate the respective simulation images.
For our experiments, we manufacture $4$ blue-coloured DLOs using Shore hardness 00-20 (for len. 200mm and 290mm), 00-50 (for len. 270mm) and A-40 (for len. 200mm) silicone polymers [33]. The real world green target ball is made of plasticine.
@@ IV-C Perception setup
We collect visual observations using the real-time RGB image stream produced from our side-view camera. For computational efficiency, we focus our observations on our blue DLO and the green target ball using segmentation images. For this, we fine-tune the segmentation task version of YOLOv8.2 [34]. We use a dataset of $183$ manually labelled workspace images, pre-processed through auto-orientation and resizing to $256\times 256$, and augmented using random Gaussian blur and salt-and-pepper noise, totalling $439$ images. With this dataset, we train the open source YOLO weights for another $64$ epochs.
We reduce the dimensionality of our visual observations by applying a learnt model of $4$ keypoints to each segmentation image in real time. We train this keypoint model using a dataset of $1500$ random policy rollouts of our simulated control task, while sampling our system parameters $\boldsymbol{\theta}$ from a uniform prior. In this data set, we also incorporate a small number of real-world policy rollouts, and we train for $50$ epochs. These keypoints constitute our observation vector.
@@ IV-D Domain Randomisation setup
Thus, to perform DR for our RL task, we train our policies in vectorised environments [32]. This enables us to launch parallel instances of our simulation, each with a different DLO parameterisation, which is sampled by our current reference prior. To further robustify the real-world performance of our sim-trained policies and minimise the need for camera calibration, we introduce a small randomness in our camera and target object poses by uniformly sampling pose offsets in the [±$[\pm 0.025,\;\pm 0.025,\;\pm 0.025]$ (m) and [±0.02, 0,±0.02]plus-or-minus0.02 0plus-or-minus0.02[$\pm$ 0.02,\;0,\;$\pm$ 0.02][ ± 0.02 , 0 , ± 0.02 ] (m) ranges, respectively.
We empirically select to launch $12$ concurrent environments, from $12$ respective domain samples, to manage our computationally intensive deformable object simulations. This places the descriptiveness of our inferred posteriors at the forefront of our experimental study, since inaccuracy and imprecision of the inferred MoGs can have worse consequences in small-data experiments. We sample our MoGs in a low-variance method, thus each component is likely to contribute to our set of domain samples.
@@ IV-E Control setup
We command our EEF by sending it target poses ($7$D command vectors). For our control experiments, we constrain our EEF motion into two dimensions, moving only along the $x$ and $z$ axes by controlling the respective deltas. Thus, our sampled policy actions are $2$D $\langle dx,dz\rangle$ vectors that we sample in the $[-0.06,0.06]$ (m) range to maintain smooth EEF transitions. Our $12$D observations are constructed by a flattened vector of the $2$D desired goal, the $2$D vector of the $x$ and $z$ components of our EEF position (proprioception) and the $4\times 2$D state observation, which reflects the pixel positions of our target object and DLO, respectively.
In each state $s_{t}$, we track the current distance of our DLO keypoint cluster to the desired goal, which we compute using the Frobenius norm of the respective distance matrix. Our sparse reward function uses this distance $d_{t}$ and a predefined pixel space distance threshold to count the reward $d_{\text{thresh}}$. Thus, whenever $d_{t}\leq d_{\text{thresh}}$, our respective reward is the distance scaled to $[0.0,1.0]$, denoted as $r_{t}=1.0-(d_{t}/d_{\text{thresh}})$.
For safety reasons, we restrict our EEF motion within the $\langle x,y,z\rangle\in\langle[0.275,0.6],[-0.1,0.1],[0.1,0.5]\rangle$ (m) world frame coordinates, with the base of the robot arm being on $(0,0,0)$. Whenever the EEF leaves this designated workspace, the episode ends as a failure with a reward of $-1$.
@@ IV-F Parameter Inference & Policy Algorithms training
For parameter inference, we implement the BayesSim-RKHS variant [14] (Fig. 2). Our RKHS-net layer embeds the $5$ observed keypoints ($4$ from DLO, $+1$ for the desired goal) using $500$ random Fourier features and cos-only components. Our MDNN models $4$ mixture components with full covariance and uses $3$ fully connected layers of $1024$ units each. We use Adam optimiser with a learning rate of $1e$$-6$. As in Alg. 1, line 8, we approximate our posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each increasing the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ V-B Uncertainty over Parameter Estimation & DR
Following the above property of MoG posteriors, we observe how the relative certainty (sharpness) of a posterior along the dimension of a variable results in proportionately tightly clustered domain samples along this dimension. Thus we see domain samples being mainly spread along the less certain – length – posterior dimension. This is particularly visible for the notably similar 200mm and 290mm shore 00-20 DLO posteriors.
For the 200mm; A-40 and 270mm; 00-50 DLOs, we see some alternative hypotheses (component means) for the DLO parameterisation that also contribute to our low-variance sampling, due to the relative magnitude of the respective mixture coefficients. This contribution is evident by the spread of the respective domain samples.
@@ V-C Object-centric Agent Performance
These mean trajectories are presented in Fig. 5, with shaded regions showing the standard deviation between different repetitions. We observe many interesting motion patterns, which indicate adaptation of agent performance to the physical parameterisation of the respective evaluation DLO. An indicative subset of these observations is extrinsically corroborated by Fig. 6. The following observations are highlighted in olive annotations.
@@
Figure 5: EEF trajectories during the real world deployment for $4$ different DLOs of $6$ policies trained using different domain distributions. We repeat each deployment $4$ times and average the measured accumulation of commanded EEF translations along the $x$ and $z$ axes (commanded actions), with shaded regions reporting standard deviation.
[…]
Figure 7: Dynamic Time Wrapping (DTW) heatmap of EEF trajectories (accum. commanded actions) following trained PPO policies.
@@ VI Discussion
In this paper, we extend the Real2Sim2Real literature with a proof of concept that an integrated distributional treatment of parameter inference, policy training, and zero-shot deployment can lead to notable agent behaviour adaptation over a set of parameterised DLOs. We use a prominent LFI method (BayesSim) to capture fine parameterisation differences of similarly shaped DLOs. Using the inferred MoG posteriors for DR during RL policy training may not lead to direct benefits in training performance (although this is largely task dependent); however, we observe strong object-centric agent performance indications during real-world deployment.
This means that through an integrated distributional treatment of the Real2Sim2Real problem, practitioners of the soft robotics and soft body manipulation domains can condition RL policies on system parameterisations inferred through visual observations, thus inducing object-centric performance, all the while operating in a shared action and observation space for both parameter inference and policy learning. Finally, we note that our framework can be extended for dual control by improving the posterior estimation in tandem with the task policy [21, 36].
ACKNOWLEDGMENTS
G. Kamaras is supported by the Engineering and Physical Sciences Research Council (EPSRC), as part of the CDT in RAS at Heriot-Watt University and The University of Edinburgh. S. Ramamoorthy is supported by a UKRI Turing AI World Leading Researcher Fellowship on AI for Person-Centred and Teachable Autonomy (grant EP/Z534833/1).

### v2

@@ IV-B Simulation & Hardware setup
We setup our simulation in IsaacGym [30]. Our simulated environment contains a Franka Emika Panda 7-DoF robot arm equipped with a parallel gripper, which is on top of a tabletop workspace. On top of this workspace, a blue DLO exists, implemented as a tetrahedral grid, and simulated using the corotational finite element method of the FleX physics engine. A small green spherical object is our task’s reaching target. Across all our parameter inference and policy learning experiments, the simulated DLO is parameterised in the $[1e3,5e4]$ (Pa) range for Young’s modulus and the $[195,305]$ (mm) range for length. Thus, median of our simulated parameter space is $\mu=(2.55e4\;\text{Pa},\;250\text{mm})$.
Our real-world setup closely replicates the simulation setup. We extend an open-source sample-efficient robotic RL infrastructure [29] for the purpose of evaluating PPO policies [31]. For visual observations, we use a RealSense D435i camera capturing $60$ fps, mounted to view our workspace from the right side. To avoid the overhead of a more elaborate sim and real camera calibration, we select our real camera placement so that the captured images qualitatively approximate the respective sim images similar to [32, 14].
[…]
The real green target ball is made of plasticine. It is simulated with a fixed rigid sphere of radius $r=1$cm. Since DLO-target physical interactions are irrelevant to our task, this reality gap of the target material is also irrelevant.
TABLE I: Real DLO indexes and parameterisations.
DLO idx. 0 1 2 3
Length 200mm 200mm 270mm 290mm
Shore A-40 00-20 00-50 00-20
Hardness (medium soft) (extra soft) (soft) (extra soft)
Mass $47.47$g $42.94$g $57.8$g $63.94$g
@@ IV-C Perception setup
We collect visual observations using the real-time RGB image stream of our side-view camera. For computational efficiency, we focus our observations on our blue DLO and the green target ball using segmentation images. For this, we fine-tune the segmentation task version of YOLOv8.2 [34]. We use a dataset of $183$ manually labelled workspace images, pre-processed through auto-orientation and resizing to $256\times 256$, and augmented using random Gaussian blur and salt-and-pepper noise, totalling $439$ images. With this dataset, we train the open source YOLO weights for another $64$ epochs.
We reduce the dimensionality of our visual observations by applying a learnt model of $4$ keypoints to each segmentation image in real time. We train this keypoint model using a dataset of $1500$ random policy rollouts of our simulated control task, while sampling our system parameters $\boldsymbol{\theta}$ from a uniform prior. In this data set, we also incorporate a small number of real-world policy rollouts, and we train for $50$ epochs. These keypoints are our DLO observation vector.
@@ IV-D Domain Randomisation setup
Thus, to perform DR for our RL task, we train our policies in vectorised environments [31] and launch parallel instances of our simulation, each with a different $\boldsymbol{\theta}$, sampled by our current reference prior $p(\boldsymbol{\theta})$. To further robustify the Sim2Real transfer of our policies and minimise the need for camera calibration, we introduce a small randomness in our camera and target object positions by uniformly sampling $(x,y,z)$ offsets in the (±$(\pm 0.025,\;\pm 0.025,\;\pm 0.025)$ (m) and (±0.02, 0,±0.02)plus-or-minus0.02 0plus-or-minus0.02($\pm$ 0.02,\;0,\;$\pm$ 0.02)( ± 0.02 , 0 , ± 0.02 ) (m) ranges, respectively.
We empirically chose to launch $12$ concurrent environments, from $12$ respective domain samples, to manage our computationally intensive deformable object simulations. This places the descriptiveness of our inferred posteriors at the forefront of our experimental study, since inaccuracy and imprecision of the inferred MoGs can have worse consequences in small-data experiments. We sample our MoGs in a low-variance method, thus each component is likely to contribute to our set of domain samples.
@@ IV-E Control setup
We command our EEF by sending it cartesian pose commands ($7$D vectors). For our control experiments, we constrain our EEF motion into two dimensions, moving only along the $x$ and $z$ axes by controlling the respective deltas. Thus, our sampled policy actions are $2$D $\langle dx,dz\rangle$ vectors that we sample in the $[-0.06,0.06]$ (m) range to maintain smooth EEF transitions. Our $12$D observations are constructed by flattening and concatenating our EEF’s $\langle x,z\rangle$ $2$D position vector (proprioception) and the $5\times 2$D visual observation, which are the positions of our DLO and target keypoints.
In each state $s_{t}$, we compute the distance $d_{t}$ of our DLO keypoint cluster to the target using the Frobenius norm of the respective distance matrix. Our sparse reward function uses this $d_{t}$ and a pixel space distance threshold $d_{\text{thresh}}=1.5$ to count the reward $r_{t}$. Thus, whenever $d_{t}\leq d_{\text{thresh}}$, our respective reward is the distance scaled to $[0.0,1.0]$, denoted as $r_{t}=1.0-(d_{t}/d_{\text{thresh}})$. The episode ends if $r_{t}\geq 0.75$.
For safety reasons, we restrict our EEF motion within the $\langle x,y,z\rangle\in\langle[0.275,0.6],[-0.1,0.1],[0.1,0.5]\rangle$ (m) world frame coordinates, with the robot arm based on $(0,0,0)$. Whenever the EEF leaves this designated workspace, the episode ends as a failure with a reward of $-1$.
@@ IV-F Parameter Inference & Policy Algorithms training
For parameter inference, we implement the BayesSim-RKHS variant [14] (Fig. 2). Our RKHS-net layer embeds the $5$ observed keypoints using $500$ random Fourier features and cos-only components. Our MDNN models $4$ mixture components with full covariance and uses $3$ fully connected layers of $1024$ units each. We use Adam optimiser with a learning rate of $1e$$-6$. As in Alg. 1, line 8, we approximate our posterior $\hat{p}(\boldsymbol{\theta})$ through $15$ iterations, each increasing the training set with $100$ more trajectories, whose parameters $\boldsymbol{\theta}$ have been sampled using the latest posterior.
@@ V-B Uncertainty over Parameter Estimation & DR
Following the above property of MoG posteriors, we observe how the relative certainty (sharpness) of a posterior along the dimension of a variable results in proportionately tightly clustered domain samples along this dimension. Thus we see domain samples being mainly spread along the less certain length dim. of posteriors. This is particularly visible for the notably similar posteriors of the very soft DLOs 1 and 3. For DLOs 0 and 2, we see that some of the alternative hypotheses (component means) for $\boldsymbol{\theta}$ contribute to our low-variance sampling, due to the relative magnitude of the respective mixture coefficients. This contribution is evident in the spread of the respective domain samples.
@@ V-C Object-centric Agent Performance
For each experiment, we report the mean trajectories resulting from the accumulation of commanded EEF position deltas (commanded actions, $\langle dx,dz\rangle$) over $4$ repetitions. These are presented in Fig. 5, with shaded regions showing the standard deviation between different repetitions. We observe many interesting motion patterns, which indicate adaptation of agent performance to the physical parameterisation of the respective evaluation DLO. An indicative subset of these observations is extrinsically corroborated by Fig. 6. The following observations are highlighted in olive annotations. Policy underlines match plot line colours.
@@
Despite these extrinsic adaptation indications, Fig. 5 averaged rewards and pixel-distance to the target show relatively similar quantitative results, both in real and sim trajectories. This suggests that adaptation occurs at a behavioural level not fully captured by our sparse reward function. This highlights the importance of trajectory-based evaluation in dynamic DLO manipulation. Due to drag, inertia, and non-trivial DLO dynamics, success may require nuanced motion patterns that are difficult to capture by scalar proximity metrics.
Figure 5: EEF trajectories during the deployment of $6$ policies trained using different domain distributions (col. 1-6) in real world (rows 1-4; $4$ real DLOs) and in simulation (row 5; median DLO). We repeat each deployment $4$ times and average the measured accumulation of commanded EEF translations along the $x$ and $z$ axes (commanded actions, $\langle dx,dz\rangle$), with shaded regions reporting standard deviation. Colorbars illustrate DLO keypoints pixel-distance to the target $d_{t}$, reversing reward function for $d_{\text{thresh}}=1.5$. The rightmost column shows the averaged rewards per timestep for the respective row’s trajectories.
Figure 6: Timelapses of real-world PPO 1 and 3 policy evaluations using DLOs 1 and 3. In conjuction with Figure 5, we extrinsically confirm reported observations (3) (same $1/3$ of trajectory), (4) (similar reaching pattern) and (9) (moving higher above the table). Dotted white and red lines visually mark the initial EEF $z$ and table surface respectively. A purple indicator on top left of each image marks the initial EEF position, with dotted purple lines marking its visual displacement. Dotted cyan lines in $t=\{0,15\}$ mark DLO body distance to target.
Figure 7: Dynamic Time Wrapping (DTW) heatmap of real and sim EEF trajectories (accum. commanded actions) over PPO-$\{\mathit{U},\mu,0,1,2,3\}$ evaluations.
@@ VI Discussion
In this paper, we extend the Real2Sim2Real literature with a proof of concept that an integrated distributional treatment of parameter inference, policy training, and zero-shot deployment can lead to significant agent behaviour adaptation over a set of parameterised DLOs. We use a prominent LFI method (BayesSim) to capture fine parameterisation differences of similarly shaped DLOs. Using the inferred MoG posteriors for DR during RL policy training may not lead to direct benefits in training performance (although this is largely task dependent); however, we observe strong object-centric agent performance indications during real-world deployment.
Through our framework practitioners of soft robotics and soft body manipulation can condition RL policies on system parameterisations inferred through vision, thus inducing object-centric performance, all the while operating in a shared action and observation space for both parameter inference and policy learning. In this context, a constraint to be addressed is that, although we reduce the reality gap in terms of realism (observations), we do not yet account for physical accuracy (true states). Thus, e.g., the inferred Young’s of a DLO may not match the real value, especially if the Real2Sim algorithm has to account for higher-order $\boldsymbol{\theta}$.